# Replacing Softmax Similarity with a Sharpened Angular Similarity: Theory and Practice of Scaling To Billion-Context Attention

### v1

@@ Abstract
Softmax Attention has a quadratic time complexity in sequence length, which becomes prohibitive to run at long contexts, even with highly optimized GPU kernels. For example, FlashAttention (an exact, GPU-optimized implementation of Softmax Attention) cannot complete a single forward–backward pass of a multi-head attention layer once the context exceeds $\sim\!4$ million tokens on an NVIDIA GH200 (96 GB). We introduce RACE Attention, a kernel-inspired alternative to Softmax Attention that is linear in sequence length and embedding dimension. RACE Attention replaces the exponential kernel with a sharpened angular (cosine) similarity, and approximates attention outputs via randomized projections and soft Locality-Sensitive Hashing (LSH). Across language modeling, masked language modeling, and text/image classification, RACE Attention matches or outperforms strong baselines while reducing wall-clock time and memory. In a controlled scale test, it processes up to 12 million tokens during a single forward-backward pass on an NVIDIA GH200 GPU and 75 million tokens on an Intel Xeon® Gold 5220R CPU—well beyond the practical limits of the current state-of-the-art attention implementations. RACE Attention thus offers a practical, theoretically grounded mechanism for outrageously long context windows on today’s hardware. We hope that it gets adopted in practice.
@@ 1 Introduction
The Transformer [31, 11] is the backbone of modern sequence modeling across language, vision [23], and speech [19]. We have seen remarkable improvements over the past few years in reasoning and understanding capabilities. Most of these are attributed to the increased parameters of the transformers along with the capability to process longer context windows than before. All this progress, however, rests on a computationally expensive primitive: Softmax Attention, whose time scales quadratically with context length. As models and contexts grow—from multi-document reasoning to long-form code, audio, and video—this quadratic barrier increasingly dictates who can train and deploy capable systems. Industrial labs mitigate the cost with large-scale distributed hardware; most practitioners cannot. There is a growing need for attention mechanisms that are accurate, fast, and memory-efficient. To highlight the limits of Softmax Attention: even with FlashAttention [10]—the state-of-the-art GPU implementation—a single forward–backward pass of a multi-head attention layer (1 batch, 4 heads, embedding dimension of 128) remains computationally and memory intensive and cannot process sequences beyond $\sim 4$ million tokens on an NVIDIA GH200 (96 GB). Clearly, to achieve an outrageously long context where the target context size is hundreds of millions of tokens or beyond, fundamental rethinking of attention will be required [2].
@@
where the softmax is applied row-wise so that attention weights are nonnegative and sum to one.
In this paper, we propose a surprising alternative to the softmax—namely, a high-degree monomial of an angular kernel based on cosine geometry:
@@ 2.2 RACE Sketch
RACE [8, 9] shows that any similarity expressible as a (non-negative) linear combination of LSH collision kernels can be sketched using ACE-style estimation [18]. It provides an unbiased estimator of kernel-density sums for LSH collision kernels and their powers. In particular, ACE/RACE estimates $\sum_{x\in D}k(x,q)^{p}$ by hashing items into counters and reading the counters addressed by the query; averaging across $L$ independent rows reduces variance.
@@ 3.1 Softmax-Like Similarities that Admit Linear-Time Estimation
In Figure 2, we show that for sufficiently large $\gamma$, the angular kernel becomes almost indistinguishable from softmax. This is expected because a high degree monomial like $x^{12}$ behaves similarly to an exponential.
@@
Figure 2: Comparison of Softmax and Angular kernels at different sharpening levels $\gamma$. As $\gamma$ (or non-linearity) increases, Angular transitions from flat similarity scores to a sharper distribution, recovering behavior similar to the exponential of the Softmax.
@@ 3.3 Theoretical Analysis of Algorithm 1
Remark (Causal masking): Our experiments in Section 4.1 employ RACE Attention with causal masking, implemented efficiently via OpenMP. See Algorithm 2 in the Appendix for the causal soft RACE Attention algorithm. However, our theoretical analysis above applies only to the non-causal setting. Extending the bias–variance guarantees of Theorem 2 to the masked case remains an open problem, as the cumulative-sum constraint interacts non-trivially with the random feature construction. A rigorous analysis of causal RACE Attention is an important direction for future work.
@@ 4 Experiments
To avoid cherry-picking and ensure comparability, we adopt the evaluation suites standard in prior efficient-attention work—Linear Attention, Linformer, and Random Feature Attention (RFA) [15, 32, 24]. Concretely, we include text classification (QNLI [27], SST-2 [28], IMDB [20], and Yahoo [37]) to probe moderate-length discriminative accuracy (as in Linformer); autoregressive language modeling to test token-level modeling (as in RFA) on WikiText-103 [22] and Penn Tree Bank (PTB) corpus [21]; masked language modeling on Tiny Stories dataset [13]; image classification (CIFAR-10 [17] and FashionMNIST [33]) to test expressivity using Vision Transformer [12] architecture (as in Linear Attention); and long-context reasoning via Long Range Arena [29] (e.g., ListOps and Text Retrieval) to stress scaling and accuracy. Together these cover four regimes—bidirectional, autoregressive, long-context, and moderate-context text/image classification. Beyond these standard benchmarks, we additionally report extreme-length scaling experiments to tens of millions of tokens. To the best of our knowledge, this is the first work to report experiments with attention spanning close to 100 million tokens. In this section, we introduce an additional hyperparameter $M$ representing the number of ensembles per head i.e., independent replications of the whole sketching scheme inside each attention head. For clarity, FlashAttention is an exact, fused-kernel implementation of Softmax Attention; we therefore use “FlashAttention” and “Softmax Attention” interchangeably when discussing accuracy and runtime.
@@
Baselines: We evaluate RACE against widely used baselines with publicly available implementations: FlashAttention [10], Linear Attention [15], Performer [6], and Linformer [32]. These span exact, kernel-linear, and low-rank approximations. All models are tuned per authors’ guidelines and trained under identical settings.
4.1 Is RACE Attention as Accurate as Tansformers?
We report text-classification accuracy in Tables 1, 4, 6, and 12; long-context (LRA) results—ListOps and Text Retrieval—in Tables 3 and 5; image-classification accuracy on CIFAR-10 in Table 9 and 10; and masked language modeling (MLM) perplexity in Tables 2 and 11. For autoregressive language modeling, RACE matches softmax-level perplexity on WikiText-103 and improves upon it on PTB (Tables 7, 8). These results indicate that RACE preserves accuracy in the overlapping regime while delivering consistent gains on long-context settings.
Unless stated otherwise, all methods use the same Transformer backbone (layers, heads, embedding dimension, dropout) and training budget. We train with identical optimizers, schedulers, and batch sizes; full hyperparameters appear in Table 13. Metrics are reported from the best-validation checkpoint. Despite the extra parameters introduced by Linformer’s lengthwise projections, it does not outperform RACE under matched training conditions. For the long-range task in Table 5, FlashAttention and Angular Attention have quadratic time complexity and are prohibitively slow; we therefore exclude them. All experiments were run on NVIDIA A100 GPUs.
@@ 4.2 Can we reach 100 million context window on popular hardware?
RACE scales up to 75 million tokens for a single forward-backward pass on CPU. By contrast, FlashAttention becomes prohibitively slow at around $\sim\!2$ million tokens due to the quadratic time scaling in sequence length $N$ (see figs. 3 and 4). It is worth noting that FlashAttention does not run out of memory on the CPU DRAM. However, the time required to complete a single forward-backward pass increases quadratically with $N$. RACE is more than $10000\times$ faster than FlashAttention at context length of $33$ million. RACE finishes comfortably under 10 seconds for a single forward-backward pass on this hardware while FlashAttention takes approximately $10^{5}$ seconds on the same hardware. RACE gets even faster with increasing context length and takes about 100 seconds to perform the same operation with $75$ million tokens. This is expected because RACE is linear and FlashAttention is quadratic in the number of tokens. The experiments also highlight that linear attentions’ approximations are not only inaccurate but also significantly slower and have large memory overheads due to large hidden constants. They run about an order of magnitude slower than RACE Attention and even go out of memory at around 33 million context length.
@@ 4.3 Right Algorithm beats Hardware Acceleration!
5 Discussion
We introduced RACE Attention, a linear-time, memory-efficient alternative to Softmax Attention that estimates a sharpened angular kernel via RACE sketches. By replacing explicit pairwise scores with bucketed aggregates, compute and activation memory scale linearly with context length $N$, and embedding dimension $d$, with constants governed by the sketch parameters $(P,L,M)$. In our rigorous stress-test, we scale at context lengths up to $\mathbf{75\,million}$ tokens on CPU and $\mathbf{12\,million}$ tokens on GPU (single forward–backward pass), a regime infeasible for Softmax Attention and other highly optimized linear attention mechanisms under comparable settings.
@@ Acknowledgments
Unless otherwise stated, we use a linear warmup–decay learning-rate schedule. Let $T$ denote the total number of optimizer updates ($T=\mathrm{epochs}\times\mathrm{len(train\_loader)}$)). The learning rate increases linearly from 0 to the base value over the first $0.01T$ updates, then decays linearly to 0 over the remaining T0.01TT-0.01T updates; the scheduler is stepped once per optimizer update. We use the AdamW optimizer with hyperparameters listed in Table 13. For learning rate $2e^{-4}$, we don’t use linear warmup or a scheduler, we simply train the model with the constant learning rate for certain number of epochs given in Table 13.
@@ 6.2 Kernel construction with the bias term
The deviation bound decomposes into a variance term and a (deterministic) bias term $\tilde{B}$. We now bound $\tilde{B}$ explicitly as a function of $\beta$ and $P$.
Lemma 7 (Bias to the $P$-powered angular kernel: explicit bound).
@@
Next, let $p:=\phi^{(\ell)}(Q_{i})$ and $q:=\phi^{(\ell)}(K_{j})$ denote the softmax feature vectors for $Q_{i}$ and $K_{j}$, and let $a:=r^{\star}(Q_{i})$, $b:=r^{\star}(K_{j})$ be their respective dominant corners. By the deterministic inequality
@@ 7 Causal Race Attention
We implemented the causal version efficiently using OpenMP-based parallelization rather than a naïve nested-loop approach. Each hash table is processed in a separate thread with its own cumulative bucket arrays, and updates are performed incrementally in a single left-to-right scan. This avoids redundant recomputation at every step using torch.cumsum() and enables CPU-level parallel execution with negligible synchronization overhead.

### v2

@@ Abstract
Softmax Attention has a quadratic time complexity in sequence length, which becomes prohibitive to run at long contexts, even with highly optimized GPU kernels. For example, FlashAttention (an exact, GPU-optimized implementation of Softmax Attention) cannot complete a single forward–backward pass of a multi-head attention layer once the context exceeds $\sim\!4$ million tokens on an NVIDIA GH200 (96 GB). We introduce RACE Attention, a kernel-inspired alternative to Softmax Attention that is linear in sequence length and embedding dimension. RACE Attention replaces the exponential kernel with a sharpened angular (cosine) similarity, and approximates attention outputs via randomized projections and soft Locality-Sensitive Hashing (LSH). Across language modeling, masked language modeling, and text/image classification, RACE Attention matches or outperforms strong baselines while reducing wall-clock time and memory. In a controlled scale test, it processes up to 12 million tokens during a single forward-backward pass on an NVIDIA GH200 GPU and 75 million tokens on an Intel Xeon® Gold 5220R CPU—well beyond the practical limits of the current state-of-the-art attention implementations. RACE Attention thus offers a practical, theoretically grounded mechanism for outrageously long context windows on today’s hardware. We hope that it gets adopted in practice. Code is open-sourced at https://github.com/sahiljoshi515/RACE_Attention.
@@ 1 Introduction
The Transformer [31, 11] is the backbone of modern sequence modeling across language, vision [23], and speech [19]. We have seen remarkable improvements over the past few years in reasoning and understanding capabilities. Most of these are attributed to the increased parameters of the transformers along with the capability to process longer context windows than before. All this progress, however, rests on a computationally expensive primitive: Softmax Attention, whose time scales quadratically with context length. As models and contexts grow—from multi-document reasoning to long-form code, audio, and video—this quadratic barrier increasingly dictates who can train and deploy capable systems. Industrial labs mitigate the cost with large-scale distributed hardware; most practitioners cannot. There is a growing need for attention mechanisms that are accurate, fast, and memory-efficient. To highlight the limits of Softmax Attention: even with FlashAttention [10]—the state-of-the-art GPU implementation—a single forward–backward pass of a multi-head attention layer (1 batch, 4 heads, embedding dimension of 128) remains computationally and memory intensive and cannot process sequences beyond $\sim 4$ million tokens on an NVIDIA GH200 (96 GB) GPU. Clearly, to achieve an outrageously long context where the target context size is hundreds of millions of tokens or beyond, fundamental rethinking of attention will be required [2].
@@
where the softmax is applied row-wise so that attention weights are nonnegative and sum to one. In this paper, we propose a surprising alternative to the softmax—namely, a high-degree monomial of an angular kernel based on cosine geometry:
@@ 2.2 RACE Sketch
RACE [8, 9] shows that any similarity expressible as a (non-negative) linear combination of LSH collision kernels can be sketched using ACE-style estimation [18]. It provides an unbiased estimator of kernel-density sums for LSH collision kernels and their powers. In particular, RACE estimates $\sum_{x\in D}k(x,q)^{p}$ by hashing items into counters and reading the counters addressed by the query; averaging across $L$ independent rows reduces variance.
@@ 3.1 Softmax-Like Similarities that Admit Linear-Time Estimation
In Figure 2, we show that for sufficiently large $\gamma$, the angular kernel becomes almost indistinguishable from softmax kernel. This is expected because a high degree monomial like $x^{12}$ behaves similarly to an exponential.
@@
Figure 2: Comparison of Softmax and Angular kernels at different sharpening levels $\gamma$. As $\gamma$ (or non-linearity) increases, Angular transitions from flat similarity scores to a sharper distribution, recovering behavior similar to the exponential in the Softmax.
@@ 3.3 Theoretical Analysis of Algorithm 1
Remark (Causal masking): Our language modeling experiments in Section 4.1 employ RACE Attention with causal masking, implemented efficiently via OpenMP for CPU workloads. See Algorithm 2 in the Appendix for the causal soft RACE Attention algorithm. Our theoretical analysis above applies only to the non-causal setting. Extending the bias–variance guarantees of Theorem 2 and creating an efficient CUDA kernel for the masked case remains an open problem, as the cumulative-sum constraint interacts non-trivially with the random feature construction. A rigorous analysis of causal RACE Attention is an important direction for future work.
@@ 4 Experiments
To avoid cherry-picking and ensure comparability, we adopt the evaluation suites standard in prior efficient-attention work—Linear Attention, Linformer, and Random Feature Attention (RFA) [15, 32, 24]. Concretely, we include text classification (QNLI [27], SST-2 [28], IMDB [20], and Yahoo [37]) to probe moderate-length discriminative accuracy (as in Linformer); autoregressive language modeling to test token-level modeling (as in RFA) on WikiText-103 [22] and Penn Tree Bank (PTB) corpus [21]; masked language modeling on Tiny Stories dataset [13]; image classification (CIFAR-10 [17] and FashionMNIST [33]) to test expressivity using Vision Transformer [12] architecture (as in Linear Attention); and long-context reasoning via Long Range Arena benchmark [29] (e.g., ListOps and Text Retrieval) to stress scaling and accuracy. Together these cover four regimes—bidirectional, autoregressive, long-context, and moderate-context text/image classification. Beyond these standard benchmarks, we additionally report extreme-length scaling experiments to tens of millions of tokens. To the best of our knowledge, this is the first work to report experiments with attention spanning close to 100 million tokens. In this section, we introduce an additional hyperparameter $M$ representing the number of ensembles per head i.e., independent replications of the whole sketching scheme inside each attention head. For clarity, FlashAttention is an exact, fused-kernel implementation of Softmax Attention; we therefore use “FlashAttention” and “Softmax Attention” interchangeably when discussing accuracy and runtime.
@@
Baselines: We compare RACE against widely used baselines with publicly available implementations: FlashAttention [10], Linear Attention [15], Performer [6], and Linformer [32]. These span exact, kernel-linear, and low-rank approximations. All models are tuned per authors’ guidelines and trained under identical settings.
4.1 Is RACE Attention as Accurate as Transformers?
We report text-classification accuracies in Tables 1, 4, 6, and 12; long-context (LRA) results—ListOps and Text Retrieval—in Tables 3 and 5; image-classification accuracies in Table 9 and 10; and masked language modeling (MLM) perplexities in Tables 2 and 11. For autoregressive language modeling, RACE matches softmax-level perplexity on WikiText-103 and improves upon it on PTB (Tables 7, 8). These results indicate that RACE preserves accuracy in the overlapping regime while delivering consistent gains on moderate-context and long-context settings.
Unless stated otherwise, all methods use the same Transformer backbone (layers, heads, embedding dimension, dropout) and training budget. We train with identical optimizers, schedulers, and batch sizes; full hyperparameters appear in Table 13. Metrics are reported from the best-validation checkpoint. Despite the extra parameters introduced by Linformer’s lengthwise projections, it does not outperform RACE under matching training conditions. For the long-range task in Table 5, FlashAttention and Angular Attention have quadratic time complexity and are prohibitively slow; we therefore exclude them. All experiments were run on NVIDIA A100 GPUs.
@@ 4.2 Can we reach 100 million context window on popular hardware?
RACE scales to 75 million tokens for a single forward-backward pass on CPU. By contrast, FlashAttention becomes prohibitively slow at $\sim\!2$ million tokens due to the quadratic time scaling in sequence length $N$ (see figs. 3 and 4). It is worth noting that FlashAttention does not run out of memory on the CPU DRAM. RACE is more than $10000\times$ faster than FlashAttention at context length of $\sim\!33$ million. RACE finishes comfortably under 10 seconds for a single forward-backward pass on this hardware while FlashAttention takes approximately $10^{5}$ seconds on the same hardware. Although the absolute runtime grows with $N$, RACE’s speedup over FlashAttention increases. At 75 million tokens, RACE finishes in about 100 seconds. This is expected because RACE is linear and FlashAttention is quadratic in $N$. The experiments also highlight that linear attentions’ approximations are not only inaccurate but also significantly slower and have large memory overheads due to large hidden constants. They run about an order of magnitude slower than RACE Attention and even go out of memory at $\sim\!33$ million tokens.
@@ 4.3 Right Algorithm beats Hardware Acceleration!
5 Conclusion
We introduced RACE Attention, a linear-time, memory-efficient alternative to Softmax Attention that estimates a sharpened angular kernel via RACE sketches. By replacing explicit pairwise scores with aggregated bucket scores, compute and activation memory scale linearly with context length $N$, and embedding dimension $d$, with constants governed by the sketch parameters $(P,L,M)$. In our rigorous stress-test, we scale at context lengths up to $\mathbf{75\,million}$ tokens on CPU and $\mathbf{12\,million}$ tokens on GPU (single forward–backward pass), a regime infeasible for Softmax Attention and other highly optimized linear attention mechanisms under comparable settings.
@@ Acknowledgments
Unless otherwise stated, we use a linear warmup–decay learning-rate schedule. Let $T$ denote the total number of optimizer updates ($T=\mathrm{epochs}\times\mathrm{len(train\_loader)}$)). The learning rate increases linearly from 0 to the base value (provided in Table 13) over the first $0.01T$ updates, then decays linearly to 0 over the remaining T0.01TT-0.01T updates; the scheduler is stepped once per optimizer update. We use the AdamW optimizer with hyperparameters listed in Table 13. For learning rate $2e^{-4}$, we don’t use linear warmup or a scheduler, we simply train the model with the constant learning rate for certain number of epochs given in Table 13.
@@ 6.2 Kernel construction with the bias term
The deviation bound decomposes into a variance term and a (deterministic) bias term $\tilde{B}$. We now bound $\tilde{B}$ explicitly as a function of $\beta$ and $P$. Before stating the result, we introduce the following technical result that proves a deterministic inequality and will be required to bound $\|\tilde{B}\|_{2}$.
Lemma 7.
Let $p,q\in\mathbb{R}^{R}$ be probability vectors with nonnegative entries summing to one. Let $a\in\arg\max_{r}p_{r}$ and $b\in\arg\max_{r}q_{r}$. Then
$\bigl|p^{\top}q-\mathbf{1}\{a=b\}\bigr|\;\leq\;(1-p_{a})+(1-q_{b}).$
Proof.
We consider two cases.
Case 1: $a=b$. Then $\mathbf{1}\{a=b\}=1$ and
$|p^{\top}q-1|=1-p^{\top}q\leq 1-p_{a}q_{a}\quad\text{since }p^{\top}q\geq p_{a}q_{a}.$
By the inequality $(1-x)(1-y)\geq 0\Rightarrow 1-xy\leq(1-x)+(1-y)$, we obtain
$1-p^{\top}q\leq(1-p_{a})+(1-q_{a})=(1-p_{a})+(1-q_{b}).$
Case 2: $a\neq b$. Then $\mathbf{1}\{a=b\}=0$. Because $b$ maximizes $q$, we have $q_{r}\leq q_{b}$ for all $r$ and $q_{a}\leq 1-q_{b}$ (since the total mass outside $b$ is $1-q_{b}$). Hence
$p^{\top}q=p_{a}q_{a}+\sum_{r\neq a}p_{r}q_{r}\leq p_{a}(1-q_{b})+q_{b}\!\!\sum_{r\neq a}p_{r}=p_{a}(1-q_{b})+(1-p_{a})q_{b}.$
Finally,
$(1-p_{a})+(1-q_{b})-\bigl[p_{a}(1-q_{b})+(1-p_{a})q_{b}\bigr]=2(1-p_{a})(1-q_{b})\geq 0,$
so $p^{\top}q\leq(1-p_{a})+(1-q_{b})$.
Combining both cases gives the desired bound. ∎
Lemma 8 (Bounding the bias term).
@@
Next, let $p:=\phi^{(\ell)}(Q_{i})$ and $q:=\phi^{(\ell)}(K_{j})$ denote the softmax feature vectors for $Q_{i}$ and $K_{j}$, and let $a:=r^{\star}(Q_{i})$, $b:=r^{\star}(K_{j})$ be their respective dominant corners. By the deterministic inequality in Lemma 7
@@ 7 Causal Race Attention
We implemented the causal version (see Algorithm 2) efficiently using OpenMP-based parallelization rather than a naive nested-loop approach. Each hash table is processed in a separate thread with its own cumulative bucket arrays, and updates are performed incrementally in a single left-to-right scan. This avoids redundant recomputation at every step using torch.cumsum() and enables CPU-level parallel execution with negligible synchronization overhead.